# Cam-2-Cam: Exploring the Design Space of Dual-Camera Interactions for Smartphone-based Augmented Reality

### v2

@@ 3.3 Data Analysis
Table 1: Survey Questions participants answered next to corresponding AR Design Heuristics for dynamic interactions from literature that informed our post-study questionnaire. In addition to this, participants were asked if they could think of application scenarios that can leverage dual-camera mobile AR and demographic questions.
4 Results
@@
4.1 Context-Aware Dual-Camera Interplay
@@
4.1.1 Contextually Aligned Front-facing Camera Input with Virtual Reactions Shown in Rear Camera:
@@
This disconnect described during the experiment aligns with the first AR Design Heuristics presented by Endsley et al. [15]. The observed disconnect is essential to note as to provide a cohesive interaction experience for a dual-camera set-up, users should feel a sense of connectedness on their side of the phone (front-facing camera) to the virtual reaction shown in the rear camera. Although, a contrasting attitude towards Mouth Craft was described by P8 where they tell their positive experience with the mismatch of input and AR scene: “So that felt more surreal, but I think that kind of makes it more fun because it’s kind of like ‘Dr. Seussy’.” In this case, the participant referenced a children’s book author, Dr. Suess, known for their otherworldly stories that aren’t always aligned with real-world semantics. Therefore, P8 makes a connection between a familiar artistic form of expression to contextualize their experience with their unique subjective interpretation.
@@
4.1.2 Alignment of Input Types and Camera Views:
@@
4.2 Reinforcing Dual-Camera Reactions
@@
4.2.1 Bodily Action and Responsiveness:
@@
Table 2: Themes and codes with Exemplar Quotes (heat map shading indicates prevalence of category)
4.2.2 Affirming Multimodal Feedback in Dual-Camera AR:
@@
4.3 Cohesive Interaction Across Cameras
@@
4.3.1 Seamlessness of camera transitions:
Uninterrupted changes between cameras or seamlessness were described by (11, 37%) participants. When asked about their feelings towards their physical gestures and virtual reactions (Q2), P14 describes an obvious cause-and-effect between the front-facing camera delegated to capture their gesture and rear-facing camera intended to display the AR reactions: “In the first one (Face TriggAR), it was pretty much seamless… When I close one eye, it does that (virtual reaction).”. P14’s sentiment seemed related and supported by P25’s response: “It felt, I felt like there was, like, a smooth transition for like, all of them, like the graphics you can see it slowly, like, appearing.”. P25 details their experience in a way similar to how seamlessness is described in camera transitions or AR/VR transitions on HMDs, where the emphasis is placed on the gradual change between states (e.g., action and reaction) [34, 27, 8, 39, 40, 45, 41, 49].
@@ 4.3.1 Seamlessness of camera transitions:
4.3.2 Visual affirmation of physics:
@@
5 Discussion
In this section, we expand on emergent themes identified in responses from our semi-structured interviews for the Cam-2-Cam interaction concept that may contribute to new design paradigms and motivate future research venues in smartphone-based AR contexts with dual-camera setups.
5.1 Design Lessons for a Cohesive Dual-Camera Interaction Space (RQ1)
We contribute to smartphone-based AR research that leverages dual-camera setups by capturing user impressions via our questionnaire aiming to identify design lessons related to cohesive interaction between the front-facing camera and the rear-facing camera. Our concept of a coherent dual-camera interaction space adapts the definition of cohesiveness we outline in Section Cam-2-Cam: Exploring the Design Space of Dual-Camera Interactions for Smartphone-based Augmented Reality to focus on designing a dual-camera interface where the spaces captured by the front and rear cameras can be perceived as connected to each another.
@@ 5.1 Design Lessons for a Cohesive Dual-Camera Interaction Space (RQ1)
5.1.1 Design Lesson – Balancing Contextual Relevance and Feedback Quality (Themes #1 and #2):
@@
Molnar et al. relate the immediacy of feedback from AR elements and the type of feedback in response to actions to flow and immersion [29] . In addition, immersion in an AR context can be defined as the presence and connection to the AR elements with which users interact [29]. The majority of positive responses for Mouth Craft were categorized under the theme, Reinforcing Dual-Camera Reactions, which implies the experience of the application can be attributed to its multimodal feedback in response to users’ gestures.
@@
P15 echoes the immediacy of the feedback as benefiting the translation between their physical gestures and virtual reaction (Q2): “It felt pretty natural and consistent — when I opened my mouth, for the most part, it would place a block.” and, “There was a connection between what I was doing with my body and what was appearing on the screen … I noticed it most with the mouth one (Mouth Craft) where I would do a thing, and then the tiny computer in my hand would respond to it by creating something I could see.”. What P15 describes is a sense of connectedness or embodiment to the virtual objects, which plays a critical role in the overall immersion of an AR interface [17, 40]. Additionally, despite the unnatural or unexpected virtual response, Mouth Craft was still able to induce a ‘natural’ feeling for P15. These sentiments could indicate a design opportunity to make more engaging smartphone AR applications by introducing a dual-camera setup where a user’s space can feel connected to the space captured by the rear-facing camera.
@@
Responsiveness and quality feedback supplementing feelings of embodiment and immersion have been widely studied in smartphone AR and other XR form factors. In smartphone AR games, Xu et al. investigate how a sense of embodiment and connections to virtual characters or AR elements can be achieved with immediate, sound, and visual responses [51, 16]. On the opposite end of the spectrum, the literature demonstrates how a lack of haptic, sound, and visual cues can cause a disconnect or reduced sense of embodiment [1]. Our work extends this prior literature by demonstrating a causal relationship between the immediacy of AR feedback and multimodal sensory cues. Further, the immediacy of feedback and multimodal cues have been shown to induce feelings of connectedness in non-AR contexts [21].
@@
We believe the dual-camera setup for smartphone AR opens up a design opportunity for making the space captured by both the front-facing and rear-facing cameras be perceived as connected, thus expanding the perceived space and possibly creating a more engaging experience previously limited by the form factor. Once the gestures captured in the users’ space feel connected to what is rendered on the other side of the phone through the rear-facing camera, the embodiment the user experiences with the virtual elements may increase their presence and overall engagement, which may lead to broader adoption of smartphone AR as being a more common medium for AR experiences. Most popular smartphone AR experiences are constrained to a touch-based interaction, which is an everyday reminder that the AR scene is on the phone, whereas in immersive AR experiences, you almost want the user to forget about the form factor involved completely [48, 36].
@@
Based on our findings, a research opportunity exists to compare single-camera touched-based interaction to the dual-camera setup for AR to identify more precise trade-offs about immersion. Creating experiences where smartphone-based interactions allow users to be more immersed could lead to broader adoption as a common theme with all XR form factors is a lack of success due to poor immersion induced by their interfaces [17, 40].
@@
5.1.2 Design Lesson – Preventing Disorientation using Simultaneous Capture and Alternating Cameras (Theme #1 & Theme #3)):
@@
Face TriggAR and Mouth Craft had strengths regarding the seamlessness of captured gestures and virtual reactions. Still, a common drawback for participants was being unable to tell if their faces were in the frame for their gestures to be captured. This disorientation could be addressed with visual cues, such as a picture-in-picture (PiP) view that displays the user’s face from the front-facing camera feed or an avatar overlaid onto the rear camera feed, similar to the ‘split screen’ view suggested by P1. Similar adaptations of this type of PiP GUI design are widely used outside of AR contexts (e.g.,video calls) for self-orientation[47].
@@
Cohesive Interaction Across Cameras captured sentiments from users related to their own body orientation with respect to the phone and virtual content which may indicate a separate cameras or scenes could be made to feel connected with the combination of transitions and visual affirmation of physics. These concepts, which are analogous to cinematic camera techniques, suggest that researchers may discover new overlaps between XR and cinematic effects, potentially enabling more effective action triggers and navigation in AR. A study focused on different visual transitions or camera responsiveness (e.g., simultaneous capture or alternation speed) could shed light on the Just Noticeable Difference (JND) for users between their physical actions and virtual transitions in a multi-camera smartphone AR setup, helping to quantify strategies for perceptibly connecting spaces. Further, dual-camera setups with dual views may introduce new interaction techniques that circumvent the issues of disorientation experienced with single-camera smartphone AR, where users typically have to hold the phone while keeping their manipulating hand in-view, thus occluding parts of their arm. These spaces partitioned by the smartphone may cause incongruence between the user’s hand captured by the rear-camera and their body. [18, 38, 26] Our work encourages future research to explore various methods of affirming user actions—such as self-viewing PiP GUIs or visual cues—that may help create cohesive virtual spaces in which users experience a sense of continuity between their actions and the XR content.
5.2 Considerations of User Defined Strategies for Dual-Camera AR (RQ2)
@@
We identified strategies users developed on their own that were not part of the application interface or the experiment instructions, and these strategies provide valuable insights when designing dual-camera AR setups on smartphones. These strategies may be helpful considerations when designing smartphone AR with dual-camera AR. For example, P10 felt that head orientation was being registered with Face TriggAR: “From what I remember, me moving my head also changed the direction of the screen. I think that implied a sense of control there–the head movement determined the way the thing is sprung out (shot out)Face TriggAR), as opposed to a static image that plays every time.” Similarly, P8 demonstrated to the interviewers how she tried to move her head to place objects for Mouth Craft in the direction she wanted to, even though she was aware she needed to move her hand holding the phone to change the direction. These responses may present a design opportunity to integrate head movements into the dual-camera interface as it may be a natural response from users. Integrating an additional parameter like head rotation may also induce feelings of engagement as more physical input, especially natural ones, has been shown to lead to more engagement in AR research [50, 36]. Prior work has explored different theories within AR contexts, like Self-Determination Theory, to support users’ agency or flexibility in how they interact with technology to provide a more natural driver for sustained engagement [10, 46].
@@
Participant responses also revealed that users created mental maps connecting their space and gestures captured by the front-facing camera to the AR scene in the rear camera’s feed. Most notably, P28, P8, and P12 describe very similar mental maps when using Mirror ThrowAR, where they would make minor adjustments throughout their throws and eventually began understanding where their projectile landing location would be. The strategy required them to adjust based on where their hand was within the bounds of the screen as described by P12: “So if a ball doesn’t, let’s say, show up, or like, appear as I expected it, I would see how my hand looks on screen to try to adjust it.”
A deeper investigation into different mental mapping strategies and how they could be used to bridge separate AR spaces could be useful for smartphone AR to make dual-camera AR practical and to make mental maps that make it easier for people to perceive the extended interaction space. These user-defined strategies also support self-viewing interfaces as this would logically reduce the mental load of creating a mental map, as mentioned in the earlier design lessons. Prior work on mirror or self-viewing interfaces has suggested a more personal, meaningful, and embedded experience in contrast to interfaces where the user cannot view themselves[23, 24, 5] since mirror interfaces can deliver higher levels of personalization and immersion [35, 43, 31]. Further, this design consideration can enable a sense of competence through immediate visual feedback of observing how users interact with virtual artifacts with respect to their body, thus making interaction and orientation easier for them [5, 54].
5.3 Application Scenarios
Figure 6: After trying out Cam-2-Cam interactions, participants shared application scenarios that they thought could greatly leverage a dual-camera setup to incorporate physical gestures.
@@ 5.3 Application Scenarios
Many participants shared application scenarios that can leverage smartphone AR with a dual-camera setup, as showcased in Fig. 6. P21 described a dual-camera AR museum application where closing either eye augments an art piece you’re viewing: “So maybe with your like, if you have two eyes open, you’re seeing version A, one eye open, you’re seeing version B, when I the other eye version C.” This augmentation could add more interactive experience for art exhibits where the subject viewing the artwork may become more engaged with the work. This application scenario could also lead to research centered on embodiment with art pieces.
@@
The application scenarios in our participants’ responses applied to different domains, such as art experiences, training, gaming and education. The variety of potential applications demonstrates an extensive design space smartphone AR with dual-camera setups may present. It is believed in scenario-based design, if there are a variety of use-cases for a particular system a multitude of interaction methods are likely possible [13, 12, 33, 20]. Our study is just a first step in exploring the dual-camera smartphone AR design space, but we foresee a diverse range of interaction styles and potential research venues that can leverage our qualitative data to inform future experiments.
6 Conclusion

### v5

@@ 4.3. Data Analysis
Table 1. Survey Questions participants answered next to corresponding AR Design Heuristics for dynamic interactions from literature that informed our post-study questionnaire. In addition to this, participants were asked if they could think of application scenarios that can leverage dual-camera mobile AR and demographic questions.
5. Results
@@
5.1. Context-Aware Dual-Camera Interplay
@@
5.1.1. Contextually Aligned Front-facing Camera Input with Virtual Reactions Shown in Rear Camera:
@@
This disconnect described during the experiment aligns with the first AR Design Heuristics presented by Endsley et al. (Endsley et al., 2017). The observed disconnect is essential to note as to provide a cohesive interaction experience for a dual-camera set-up, users should feel a sense of connectedness on their side of the phone (front-facing camera) to the virtual reaction shown in the rear camera. Although, a contrasting attitude towards Mouth Craft was described by P8 where they tell their positive experience with the mismatch of input and AR scene: “So that felt more surreal, but I think that kind of makes it more fun because it’s kind of like ‘Dr. Seussy’.” In this case, the participant referenced a children’s book author, Dr. Suess, known for their otherworldly stories that aren’t always aligned with real-world semantics. Therefore, P8 makes a connection between a familiar artistic form of expression to contextualize their experience with their unique subjective interpretation.
@@
5.1.2. Alignment of Input Types and Camera Views:
@@
5.2. Reinforcing Dual-Camera Reactions
@@
5.2.1. Bodily Action and Responsiveness:
@@
Table 2. Themes and codes with Exemplar Quotes (heat map shading indicates prevalence of category)
5.2.2. Affirming Multimodal Feedback in Dual-Camera AR:
@@
5.3. Cohesive Interaction Across Cameras
@@
5.3.1. Seamlessness of camera transitions:
Uninterrupted changes between cameras or seamlessness were described by (11, 37%) participants. When asked about their feelings towards their physical gestures and virtual reactions (Q2), P14 describes an obvious cause-and-effect between the front-facing camera delegated to capture their gesture and rear-facing camera intended to display the AR reactions: “In the first one (Face TriggAR), it was pretty much seamless… When I close one eye, it does that (virtual reaction).”. P14’s sentiment seemed related and supported by P25’s response: “It felt, I felt like there was, like, a smooth transition for like, all of them, like the graphics you can see it slowly, like, appearing.”. P25 details their experience in a way similar to how seamlessness is described in camera transitions or AR/VR transitions on HMDs, where the emphasis is placed on the gradual change between states (e.g., action and reaction) (Pointecker et al., 2022; Mayer et al., 2024; Auda et al., 2023; Roo and Hachet, 2017; Sakashita et al., 2023; Tseng and Tseng, 2013; Shan et al., 2024).
@@ 5.3.1. Seamlessness of camera transitions:
5.3.2. Visual affirmation of physics:
@@
6. Discussion
6.1. Design Lessons for a Cohesive Dual-Camera Interaction Space (RQ1)
We contribute to smartphone-based AR research that leverages dual-camera setups by capturing user impressions via our questionnaire aiming to identify design lessons related to cohesive interaction between the front-facing camera and the rear-facing camera. Our concept of a coherent dual-camera interaction space adapts the definition of cohesiveness we outline in Section 1 to focus on designing a dual-camera interface where the spaces captured by the front and rear cameras can be perceived as connected to each another.
@@ 6.1. Design Lessons for a Cohesive Dual-Camera Interaction Space (RQ1)
6.1.1. Design Lesson – Balancing Contextual Relevance and Feedback Quality (Themes #1 and #2):
@@
Molnar et al. relate the immediacy of feedback from AR elements and the type of feedback in response to actions to flow and immersion (Molnár and Szűts, 2019) . In addition, immersion in an AR context can be defined as the presence and connection to the AR elements with which users interact (Molnár and Szűts, 2019). The majority of positive responses for Mouth Craft were categorized under the theme, Reinforcing Dual-Camera Reactions, which implies the experience of the application can be attributed to its multimodal feedback in response to users’ gestures.
@@
P15 echoes the immediacy of the feedback as benefiting the translation between their physical gestures and virtual reaction (Q2): “It felt pretty natural and consistent — when I opened my mouth, for the most part, it would place a block.” and, “There was a connection between what I was doing with my body and what was appearing on the screen … I noticed it most with the mouth one (Mouth Craft) where I would do a thing, and then the tiny computer in my hand would respond to it by creating something I could see.”. What P15 describes is a sense of connectedness or embodiment to the virtual objects, which plays a critical role in the overall immersion of an AR interface (Genay et al., 2021; Sakashita et al., 2023). Additionally, despite the unnatural or unexpected virtual response, Mouth Craft was still able to induce a ‘natural’ feeling for P15. These sentiments could indicate a design opportunity to make more engaging smartphone AR applications by introducing a dual-camera setup where a user’s space can feel connected to the space captured by the rear-facing camera.
@@
Responsiveness and quality feedback supplementing feelings of embodiment and immersion have been widely studied in smartphone AR and other XR form factors. In smartphone AR games, Xu et al. investigate how a sense of embodiment and connections to virtual characters or AR elements can be achieved with immediate, sound, and visual responses (Xu et al., 2011; Feuchtner and Müller, 2017). On the opposite end of the spectrum, the literature demonstrates how a lack of haptic, sound, and visual cues can cause a disconnect or reduced sense of embodiment (Abtahi et al., 2022). Our work extends this prior literature by demonstrating a causal relationship between the immediacy of AR feedback and multimodal sensory cues. Further, the immediacy of feedback and multimodal cues have been shown to induce feelings of connectedness in non-AR contexts (Iftikhar et al., 2023).
@@
We believe the dual-camera setup for smartphone AR opens up a design opportunity for making the space captured by both the front-facing and rear-facing cameras be perceived as connected, thus expanding the perceived space and possibly creating a more engaging experience previously limited by the form factor. Once the gestures captured in the users’ space feel connected to what is rendered on the other side of the phone through the rear-facing camera, the embodiment the user experiences with the virtual elements may increase their presence and overall engagement, which may lead to broader adoption of smartphone AR as being a more common medium for AR experiences. Most popular smartphone AR experiences are constrained to a touch-based interaction, which is an everyday reminder that the AR scene is on the phone, whereas in immersive AR experiences, you almost want the user to forget about the form factor involved completely (Weber et al., 2021; Qian et al., 2019).
@@
Based on our findings, a research opportunity exists to compare single-camera touched-based interaction to the dual-camera setup for AR to identify more precise trade-offs about immersion. Creating experiences where smartphone-based interactions allow users to be more immersed could lead to broader adoption as a common theme with all XR form factors is a lack of success due to poor immersion induced by their interfaces (Genay et al., 2021; Sakashita et al., 2023).
@@
6.1.2. Design Lesson – Preventing Disorientation using Simultaneous Capture and Alternating Cameras (Theme #1 & Theme #3)):
@@
Face TriggAR and Mouth Craft had strengths regarding the seamlessness of captured gestures and virtual reactions. Still, a common drawback for participants was being unable to tell if their faces were in the frame for their gestures to be captured. This disorientation could be addressed with visual cues, such as a picture-in-picture (PiP) view that displays the user’s face from the front-facing camera feed or an avatar overlaid onto the rear camera feed, similar to the ‘split screen’ view suggested by P1. Similar adaptations of this type of PiP GUI design are widely used outside of AR contexts (e.g.,video calls) for self-orientation(Vanukuru et al., 2023).
@@
Cohesive Interaction Across Cameras captured sentiments from users related to their own body orientation with respect to the phone and virtual content which may indicate a separate cameras or scenes could be made to feel connected with the combination of transitions and visual affirmation of physics. These concepts, which are analogous to cinematic camera techniques, suggest that researchers may discover new overlaps between XR and cinematic effects, potentially enabling more effective action triggers and navigation in AR. A study focused on different visual transitions or camera responsiveness (e.g., simultaneous capture or alternation speed) could shed light on the Just Noticeable Difference (JND) for users between their physical actions and virtual transitions in a multi-camera smartphone AR setup, helping to quantify strategies for perceptibly connecting spaces. Further, dual-camera setups with dual views may introduce new interaction techniques that circumvent the issues of disorientation experienced with single-camera smartphone AR, where users typically have to hold the phone while keeping their manipulating hand in-view, thus occluding parts of their arm. These spaces partitioned by the smartphone may cause incongruence between the user’s hand captured by the rear-camera and their body. (Grubert et al., 2016; Qian et al., 2020b; Ma et al., 2023) Our work encourages future research to explore various methods of affirming user actions—such as self-viewing PiP GUIs or visual cues—that may help create cohesive virtual spaces in which users experience a sense of continuity between their actions and the XR content.
6.2. Considerations of User Defined Strategies for Dual-Camera AR (RQ2)
@@
We identified strategies users developed on their own that were not part of the application interface or the experiment instructions, and these strategies provide valuable insights when designing dual-camera AR setups on smartphones. These strategies may be helpful considerations when designing smartphone AR with dual-camera AR. For example, P10 felt that head orientation was being registered with Face TriggAR: “From what I remember, me moving my head also changed the direction of the screen. I think that implied a sense of control there–the head movement determined the way the thing is sprung out (shot out)Face TriggAR), as opposed to a static image that plays every time.” Similarly, P8 demonstrated to the interviewers how she tried to move her head to place objects for Mouth Craft in the direction she wanted to, even though she was aware she needed to move her hand holding the phone to change the direction. These responses may present a design opportunity to integrate head movements into the dual-camera interface as it may be a natural response from users. Integrating an additional parameter like head rotation may also induce feelings of engagement as more physical input, especially natural ones, has been shown to lead to more engagement in AR research (Wu and Li, 2024; Qian et al., 2019). Prior work has explored different theories within AR contexts, like Self-Determination Theory, to support users’ agency or flexibility in how they interact with technology to provide a more natural driver for sustained engagement (Bennett et al., 2023; Tyack and Wyeth, 2017).
@@
Participant responses also revealed that users created mental maps connecting their space and gestures captured by the front-facing camera to the AR scene in the rear camera’s feed. Most notably, P28, P8, and P12 describe very similar mental maps when using Mirror ThrowAR, where they would make minor adjustments throughout their throws and eventually began understanding where their projectile landing location would be. The strategy required them to adjust based on where their hand was within the bounds of the screen as described by P12: “So if a ball doesn’t, let’s say, show up, or like, appear as I expected it, I would see how my hand looks on screen to try to adjust it.” Design approaches involving reducing the mental load with additional multimodal cues or a PiP view to simplify a dual-camera interface could also be considered. Such cues have suggested benefits in AR and GUIs that span spatially, allowing for more intuitive controls and enhancing feelings of embodiment (Latoschik et al., 2016; Li et al., 2021; Anderson et al., 2013).
6.3. Application Scenarios
@@
Many participants shared application scenarios that can leverage smartphone AR with a dual-camera setup as showcased in figure 1 in our appendix demonstrating breadth of potential applications. For instance, P21 described a dual-camera AR museum application where closing either eye augments an art piece you’re viewing: “So maybe with your like, if you have two eyes open, you’re seeing version A, one eye open, you’re seeing version B, when I the other eye version C.” This augmentation could add more interactive experience for art exhibits where the subject viewing the artwork may become more engaged with the work. This application scenario could also lead to research centered on embodiment with art pieces.
@@
The application scenarios in our participants’ responses applied to different domains, such as art experiences, training, gaming and education. The variety of potential applications demonstrates an extensive design space smartphone AR with dual-camera setups may present. It is believed in scenario-based design, if there are a variety of use-cases for a particular system a multitude of interaction methods are likely possible (Carroll, 2003; Carrol, 1999; O’Hare et al., 2020; Iftikhar et al., 2021). Our study is just a first step in exploring the dual-camera smartphone AR design space, but we foresee a diverse range of interaction styles and potential research venues that can leverage our qualitative data to inform future experiments.
7. Conclusion